# Learned Indexes with Distribution Smoothing via Virtual Points

### v1

@@ V CDF Smoothing for Hierarchical Indexes
The core idea is to start from the bottom most level of the index that contain parent nodes of leaf-nodes and select those nodes. Then for each of these parents nodes, the keys in the node and its child nodes are collected, which are then subjected to smoothing using Algorithm 1. If the minimum cost threshold is satisfied (more details below), then the sub-tree and the node are reconstructed to merge the collected nodes. The merging is performed by creating a new leaf-node in place of the parent node and placing the keys from the collected nodes. By doing so, more keys would be placed in upper level nodes of the index as the indexing functions of these nodes would be improved by the CDF smoothing, but the cost models would limit the number of keys as to not offset the performance gain by the increase in the leaf-node search time. Further details regarding this is given in Section V-1. This process is performed until the root node depicted as level 1 is reached, thus reducing the total prediction error ($L$). For this purpose, unbalanced learned index structures are better suited as it gives the ability to reduce the height of taller branches without affecting the rest.
@@
The algorithm for an unbalanced learned index structure is given in Algorithm 2 and described below. First the algorithm starting from the maximum level of the index, and identifies all nodes with sub-trees, which is shown in Lines 5-7. Afterwards, as shown in Lines 8-14, each of the node’s and its sub-tree’s keys are collected and subjected to the CDF smoothing. Provided that they meet the minimum cost threshold selected, the node and its sub-tree are reconstructed to promote as many keys to upper levels as possible. This process is iteratively performed in a bottom up manner for other levels of the index.
@@
V-1 Cost conditions
$\begin{split}cost=search\_constant\times expected\_number\_of\_searches\\ +traversal\_constant\times index\_level\end{split}$ (18)
To make the implementation hardware independent, the constants can be measured by sampling queries to measure the time spent per leaf-node search for the case of $search\_constant$ and the traversal time spent per level for $traversal\_constant$. The $expected\_number\_of\_searches$ can be calculated via the inbuilt function in ALEX that uses the $log_{2}$ error to estimate it. Considering the cost model depicts the expected query time for the node, the cost threshold, $c$ should be set below 0 to identify an improvement. Setting a further lower value would result in fewer keys being able to be promoted to upper levels but the expected query time improvement will be greater.
V-A Approximation Quality Analysis
@@
The key set of 10 keys given in Fig. 2 was subjected to CDF smoothing with a smoothing threshold ($\alpha$) of 0.5 (smoothing budget of 5) via both methods. The results are shown in Table II. Here, the greedy method improves the loss by 72.34%, while the exhaustive method improves it by 74.44%. However, the time taken by the exhaustive method is nearly 3 orders of magnitude more than the greedy method. This results show that the effectiveness of the greedy method is similar to the exhaustive method, and the exhaustive method is impractical to use in real datasets.
TABLE II: Approximation Quality Results
@@
VI Experimental results
Next, we report experimental results. The implementation is based on an existing benchmark [30] for easier comparison of existing techniques. All experiments were run on an Ubuntu 20.04.5 virtual machine with an AMD EPYC 7763 64-Core Processor CPU and 128 GB of RAM.
VI-A Experimental Settings
Competitors. To show the general applicability of our proposed techniques, we integrate CSV with recent learned indexes, including ALEX [2], LIPP [4], and SALI [14] (SOTA). These indexes have reported strong empirical performance, outperforming both traditional indexes like the B+-tree and learned ones like the PGM index [3], XIndex [13], and FINEdex [12]. For simplicity, we do not repeat the comparison results with these other indexes.
Datasets. We run experiments with four datasets from two benchmark works [31, 22]: (1) Facebook contains 200 million integer Facebook user IDs [32]; (2) Covid contains 200 million integer tweet IDs randomly sampled from tweets tagged with “Covid-19” [33]; (3) OSM contains 200 million locations randomly sampled from OpenStreetMap and represented as Google S2 [34] cell IDs [35]; and (4) Genome contains 200 million entries of loci pairs in human chromosomes represented as integers [36]. For all datasets, duplicate keys were removed because LIPP and SALI do not support them.
Out of the four datasets, OSM and Genome are considered more difficult for learned indexes [22] (hard datasets), while Facebook and Covid are easier (easy datasets). To illustrate this, the CDF of the full datasets are plotted in Figs. 5(a) to 5(d). All datasets except OSM have almost globally linear CDFs. Zooming in the CDFs shows that there is more variability in the local distribution patterns, as shown in Figs. 5(e) to 5(h) (each showing from the 100 million-th data point to the next thousand data points). Except Covid, all datasets deviate from linear CDFs at local level, especially Genome.
@@ VI-A Experimental Settings
Figure 5: CDFs of the datasets
@@
(1) Read-only workload. The learned indexes ALEX, LIPP, and SALI (same below) are constructed over the full datasets. Afterwards, our CSV algorithm is applied to optimise their structures. Then, the queries (detailed below) are run.
(2) Read-write workload. The learned indexes are constructed and CSV is applied over a random half of each dataset. The other halves are inserted in random insert batches of size $10\%n$. Queries are run after each batch insertion.
We consider three sets of queries over each dataset:
(1) Promoted data. Querying every key that has been promoted to upper levels in the index by our algorithm.
(2) Random. Querying one million keys randomly sampled from each dataset.
(3) Zipfian. Querying one million keys sampled from each dataset using a Zipfian distribution to represent a realistic workload, following a previous work [3].
Parameters. We vary the smoothing threshold, $\alpha$ from 0.05 to 0.8, with a default value of 0.1. To show the scalability of our algorithm by varying the dataset size, the original datasets were down sampled by eliminating every $j$-th key from the sorted datasets in order to remove $n/j$ data points and create smaller datasets of size 12.5 million, 25 million, 50 million, and 100 million, respectively. The default datasets are the original ones with 200 million data points.
For each queried key, the query time was recorded by repeating the query 100 times and taking its average. Since our main goal is to show the query performance gain achievable by our CSV algorithm, we report performance results relative to those produced by the original learned indexes without CSV, in addition to reporting the absolute query time results for random and Zipfian queries.
For LIPP and SALI, they can create nodes that are indexing only a few keys [4]. For these two indexes, CSV is run starting at the second level of the index structures, such that each smoothing step can benefit more points. This is not an issue for ALEX, and CSV is run starting at the bottom level of the structure. Further, since the query times of the keys in the top two levels of the index structures are very close, our CSV algorithm stops at the second level from the top (i.e., the root).
VI-B Results on Read-only Workloads
VI-B1 Impact of Smoothing Threshold
@@
Query time (promoted data). The total time saved by CSV, compared to the original index structure are shown in Figures 6a, 6c, and 6e for LIPP, SALI, and ALEX indexes respectively. The general trend is that adding more virtual points (i.e., increasing the smoothing budget, $\alpha$) saves more query times. LIPP and SALI tend to perform quite similarly due to SALI using LIPP as the base index. For LIPP and SALI indexes, the easy to learn datasets (Facebook and Covid) stabilise after a certain number of virtual points are inserted. This is primarily due to the original dataset’s CDF being linear. However, this is not the case for ALEX index primarily due to the leaf-node search term in its query performance.
(a) LIPP: Total time saved
(b) LIPP: Relative query time
(c) SALI: Total time saved
(d) SALI: Relative query time
(e) ALEX: Total time saved
(f) ALEX: Relative query time
Figure 6: Query time vs. smoothing threshold ($\alpha$)
(a) Percentage of
promoted data
@@
model reduction
(d) Total time saved vs.
dataset cardinality
Figure 7: (a)-(c): LIPP space cost vs. smoothing threshold, (d): LIPP query time vs. dataset cardinality
(a) Percentage of
promoted data
(b) Percentage of storage
space increase
(c) Percentage of index
model reduction
(d) Total time saved vs.
dataset cardinality
Figure 8: (a)-(c): SALI space cost vs. smoothing threshold; (d): SALI query time vs. dataset cardinality
(a) Percentage of
promoted data
(b) Percentage of storage
space increase
(c) Percentage of index
model reduction
(d) Total time saved vs.
dataset cardinality
Figure 9: (a)-(c): ALEX space cost vs. smoothing threshold; (d): ALEX query time vs. dataset cardinality
(a) LIPP: Absolute average query time (random)
(b) LIPP: Absolute average query time (Zipfian)
(c) SALI: Absolute average query time (random)
(d) SALI: Absolute average query time (Zipfian)
(e) ALEX: Absolute average query time (random)
(f) ALEX: Absolute average query time (Zipfian)
Figure 10: Query time vs. smoothing budget on other query distributions
Figs. 6(b)6(d), and 6(f) illustrate the query times of the promoted data relative to the average query times of LIPP, SALI, and ALEX, respectively. Here, the line at the $y$-value 0 represents the average query time (that is, the average query time for randomly queried data) before applying CSV (note that, we have also presented the absolute average query time for randomly queried data later in Fig. 10, where $\alpha=0.0$ represent the values for the original indexes). The dashed bars (denoted as ‘Original’) represent the query times as a percentage of the average query time for the promoted data before applying CSV. These figures show that the keys that will be promoted by CSV take a longer query time compared to the average – generally being between 15% to 80% higher than randomly queried data for LIPP and SALI. This is because the keys that will be promoted by CSV are typically at the deeper levels of the index structures that require higher query times.
After applying CSV, the query times of the promoted data as a percentage of the average query time is significantly reduced, They can be even lower than the average query time for larger values of $\alpha$ (shown by the bars below the line with $y$-value 0). As ALEX has an additional leaf-node search step which is not required by LIPP and SALI, the benefit for the promoted data is smaller (since CSV forms larger nodes that could lead to longer leaf-node search times). However, it is important to note that CSV still yields consistent query time improvements for the promoted data (i.e., the bars with solid colours are below the dashed bars of the same colour).
@@
Size of the promoted data. Figs. 7(a)8(a), and 9(a) show the number of keys promoted to higher levels after applying our CSV algorithm, as a percentage of the total number of keys that can be promoted (i.e., keys at level 3 or below). For the Facebook dataset, CSV can promote around 60% of all possible data to higher levels, while for the Covid dataset, CSV promotes around 30% of the data. For the harder to learn datasets, OSM and Genome, CSV also manages to promote up to 27% and 57% of the data, respectively. The datasets with most promoted data are again different for ALEX, which is consistent with the observations before. Also, while the general patterns are similar to those in the query times saved, higher bars may be observed on some datasets (e.g., Facebook) in these figures than those in the total time saved Figures above. This is because the promoted data is given as a percentage to the keys that can be promoted. Even though OSM and Genome report lower percentages of promoted data, the actual number of promoted data is higher. This is because, they have much higher number of keys in lower levels.
@@
Index size. Due to the addition of virtual points, we expect the storage consumption to increase. This is shown in Figs. 7(b)8(b), and 9(b). In all cases, less than 31% of additional storage space is required by the CSV-enhanced indexes compared to the original structures, and in most cases, less than 10%. The increase in the space cost is proportional to the smoothing threshold, which is also intuitive.
@@
The storage space increase is balanced out by the removal of unnecessary nodes, whose data is promoted to higher levels. This node reduction follows a similar pattern to the percentage of promoted data, as shown by Figs. 7(c)8(c) and 9(c). Here, the node reduction is given as a percentage of the nodes that could be removed (nodes at levels 3 or lower of the index structures).
Query time (other query distributions). The average query time for random distribution and Zipfian distribution queries are given in Figs. 10(a)10(c), and 10(e) and Figs. 10(b)10(d) and 10(f), respectively. When considering random queries, with the original index performance shown in $\alpha$ value of 0, their performance is slightly improved in most of the cases tested. This can be attributed to the improvements from our smoothing process. The small increase is due to a large number of keys being in higher levels that are not considered for promotion, while the random queries are mostly from those levels. The performance of the queries from the Zipfian distribution shows a similar pattern for the same reason.
@@
Pre-processing time for CSV. The times taken to run CSV to optimise the learned index structures are summarized in Tables III and IV for LIPP and ALEX. As SALI is based on LIPP, where CSV reports a similar performance, we omit the detailed results for brevity.
We see that CSV takes more time to run as the smoothing budget grows, since there are more candidate virtual points to be examined. The algorithm running times vary across different datasets, again because the datasets have different difficulties in index learning. Note that, while the running times of CSV may seem quite large under certain settings, e.g., over 80,000 seconds on OSM for ALEX, however considering CSV can be performed at the same time when the index is constructed prior to deployment, this pre-processing time is amortized by the improvement in query time.
TABLE III: CSV Pre-processing Time (s) for LIPP
0.05 0.1 0.2 0.4 0.8
@@
TABLE IV: CSV Pre-processing Time (s) for ALEX
0.05 0.1 0.2 0.4 0.8
@@
VI-B2 Impact of Dataset Cardinality.
To demonstrate the scalability of CSV against the dataset cardinality, we repeat the experiments on datasets of 12.5 million to 200 million data points. Figs. 7(d)8(d), and 9(d) show the total query times saved by applying CSV. It can be seen that for all datasets, the times saved grow with the dataset cardinality. The rate of improvement on the easier to learn datasets grows faster, since there are not many keys in the deeper levels for these datasets when the dataset cardinality is small. These results confirm the scalability of CSV towards datasets of growing cardinality.
VI-C Results on Read-write Workloads
Due to the similar trends between LIPP and SALI indexes, SALI is omitted from the results below.
Query time (promoted data). Fig. 11 shows the total query times saved by CSV, compared to the original index structures, as more batches of data are inserted. Here, the query times saved are decreasing slightly as more data is inserted for LIPP, because the inserted data have higher chance of colliding with the promoted data as they are now in higher levels, compared to when they are in lower levels as in the original index structure. For ALEX, the trend is quite similar except for on the OSM dataset, where there are two drops after one insertion batch (i.e., $0.1n$ points are inserted) and three insertion batches ((i.e., $0.3n$ points are inserted). This is primarily due to the original index structure’s query times happen to be slightly lower in these two cases.
@@ VI-C Results on Read-write Workloads
(b) ALEX
Figure 11: Total time saved vs. insertion batches (promoted data)

### v2

@@ 5. CDF Smoothing for Hierarchical Indexes
The core idea is to start from the bottom most level of the index that contains parent nodes of leaf-nodes and select those nodes. Then for each of these parents nodes, the keys in the node and its child nodes are collected, which are then subjected to smoothing using Algorithm 1. If the minimum cost threshold is satisfied (more details below), then the sub-tree and the node are reconstructed to merge the collected nodes. The merging is performed by creating a new leaf-node in place of the parent node and placing the keys from the collected nodes. By doing so, more keys would be placed in upper level nodes of the index as the indexing functions of these nodes would be improved by the CDF smoothing, but the cost models would limit the number of keys as to not offset the performance gain by the increase in the leaf-node search time. Further details regarding this is given in Section 5.1. This process is performed until the root node depicted as level 1 is reached, thus reducing the loss ($\mathcal{L}$).
This process is applied to a constructed learned index structure as the purpose of the method is to enhance the structure of the index. Further, it is computationally expensive to perform the smoothing operation on the full key set. As such it is more reasonable to handle subsets of the key set in the constructed hierarchical index. For this purpose, unbalanced learned index structures are better suited as it gives the ability to reduce the height of taller branches without affecting the rest.
@@
The algorithm for CDF smoothing of a hierarchical learned index structure is given in Algorithm 2 and described below. First the algorithm starting from the maximum level of the index, and identifies all nodes with sub-trees, which is shown in Lines 5-7. Afterwards, as shown in Lines 8-14, each of the node’s and its sub-tree’s keys are collected and subjected to the CDF smoothing. Provided that they meet the minimum cost threshold selected, the node and its sub-tree are reconstructed to promote as many keys to upper levels as possible. This process is iteratively performed in a bottom up manner for other levels of the index.
5.1. Cost Conditions
(22) $\begin{split}cost=search\_constant\times expected\_number\_of\_searches\\ +traversal\_constant\times index\_level\end{split}$
To make the implementation hardware independent, the constants can be measured by sampling queries to measure the time spent per leaf-node search for the case of $search\_constant$ and the traversal time spent per level for $traversal\_constant$. The $expected\\ \_number\_of\_searches$ can be calculated via the inbuilt function in ALEX that uses the $log_{2}$ error to estimate it. Considering the cost model depicts the expected query time for the node, the cost threshold, $c$ should be set below 0 to identify an improvement. Setting a lower value would result in fewer keys being able to be promoted to upper levels but the expected query time improvement will be greater.
@@ 5.1. Cost Conditions
Choice of smoothing threshold. Increasing the smoothing threshold would make the algorithm insert more virtual points, thus reducing the loss function value of the newly fitted model even more. As a result these indexing functions would be able to accommodate more keys which were originally in lower levels of the index and improve their query time. However, that is a trade-off between the query time improvement and the higher space cost for increasing this threshold. Further, key sets that cause the original index structure to construct poorly should benefit more from a higher smoothing threshold as shown in the Experimental results 6.2.1.
5.2. Approximation Quality Analysis
@@
The key set of 10 keys given in Fig. 2 was subjected to CDF smoothing with a smoothing threshold ($\alpha$) of 0.5 (smoothing budget of 5) via both methods. The results are shown in Table 2. Here, the greedy method improves the loss by 72.34%, while the exhaustive method improves it by 74.44%. However, the time taken by the exhaustive method is nearly 3 orders of magnitude more than the greedy method. This results show that the effectiveness of the greedy method is similar to the exhaustive method, and the exhaustive method is impractical to use in real datasets.
Table 2. Approximation Quality Results
@@
6. Experimental results
Next, we report experimental results. The implementation of our evaluation framework is based on an existing benchmark (Bachfischer et al., 2022) implementation. All experiments were run on an Ubuntu 20.04.5 virtual machine with an AMD EPYC 7763 64-Core CPU and 128 GB RAM.
6.1. Experimental Settings
Competitors. To show the general applicability of our proposed techniques, we integrate CSV with recent learned indexes, including ALEX (Ding et al., 2020a), LIPP (Wu et al., 2021), and SALI (Ge et al., 2023b) (SOTA). These three indexes were chosen because they are the latest and among the most widely used benchmark learned indexes. They have reported strong empirical performance, outperforming both traditional indexes such as the B+-tree and learned ones (Sun et al., 2023) such as the PGM index (Ferragina and Vinciguerra, 2020), XIndex (Tang et al., 2020), and FINEdex (Li et al., 2021b). For simplicity, we do not repeat the comparison results with these other indexes.
Datasets. We run experiments with four datasets from two benchmark works (Marcus et al., 2020; Wongkham et al., 2022): (1) Facebook contains 200 million integer Facebook user IDs (Van Sandt et al., 2019); (2) Covid contains 200 million integer tweet IDs randomly sampled from tweets tagged with “Covid-19” (Lopez and Gallemore, 2021); (3) OSM contains 200 million locations randomly sampled from OpenStreetMap and represented as Google S2 (S2Geometry, 2024) cell IDs (Pandey et al., 2018); and (4) Genome contains 200 million entries of loci pairs in human chromosomes represented as integers (Rao et al., 2014). For all datasets, duplicate keys were removed to suit LIPP and SALI’s requirements.
Out of the four datasets, OSM and Genome are considered more difficult for learned indexes (Wongkham et al., 2022) (hard datasets), while Facebook and Covid are easier (easy datasets). To illustrate this, the CDFs of the full datasets are plotted in Figs. 5(a) to 5(d). All datasets except OSM have almost globally linear CDFs. Zooming in the CDFs shows that there is more variability in the local distribution patterns, as shown in Figs. 5(e) to 5(h) (each showing from the 100 million-th data point to the next thousand data points). Except for Covid, all datasets deviate from linear CDFs at local level, especially Genome.
@@ 6.1. Experimental Settings
Figure 5. CDFs of the datasets
@@
(1) Read-only workload. The learned indexes ALEX, LIPP, and SALI are constructed over the full datasets. Afterwards, our CSV algorithm is applied to optimise their structures. Then, the queries (detailed below) are run.
(2) Read-write workload. The learned indexes are constructed over a random half of each dataset and then CSV is applied. The other half of the dataset is inserted in random batches of size $0.1n$. Queries are run after each batch insertion without using CSV again for each batch.
Queries. Considering the main objective of the developed method is to improve the performance of keys in lower levels, the experiments are focused on them. Specifically, we report results for the promoted data, which includes every key that has been promoted to upper levels in the index by our algorithm.
Parameters. We vary the smoothing threshold, $\alpha$, from 0.05 to 0.8, with a default value of 0.1. To show the scalability of our algorithm by varying the dataset size, the original datasets were down sampled by eliminating every $j$-th key from the sorted datasets in order to remove $n/j$ data points and create smaller datasets of size 12.5 million, 25 million, 50 million, and 100 million, respectively. The default datasets are the original ones with 200 million points.
For each queried key, the query time was recorded by repeating the query 100 times and taking its average.
For LIPP and SALI, they can create nodes that are indexing only a few keys (Wu et al., 2021). For these two indexes, CSV is run starting at the second level of the index structures, such that each smoothing step can benefit more points. This is not an issue for ALEX, and CSV is run starting at the bottom level. Further, since the query times of the keys in the top two levels of the index structures are very close, CSV stops at the second level from the top (i.e., the root).
(a) LIPP
(b) SALI
(c) ALEX
Figure 6. Total time saved vs. smoothing threshold $\alpha$
(a) LIPP
(b) SALI
(c) ALEX
Figure 7. Query time improvement vs. smoothing threshold $\alpha$
Evaluation metrics. We report: (1) the total query times saved by the CSV-enhanced indexed compared with those of the original indexes; (2) the query time improvement (%), which is the relative difference between the average query time (over all queried keys) of the CSV-enhanced indexes and that of the original ones; (3) the promoted data (%), which is the percentage of keys promoted to upper levels in the index structure among all keys that can be promoted (i.e., keys at levels 3 or below of the original indexes); (4) the storage space increase (%), which is the relative index size overhead of the CSV-enhanced indexes compared with the original ones; (5) the node reduction (%), which is the percentage of nodes reduced by the CSV-enhanced indexes over the original ones; and (6) the insert time increase (%), which is relative increase in the average time per insertion required by the CSV-enhanced indexes compared to the original ones.
6.2. Results on Read-only Workloads
6.2.1. Impact of Smoothing Threshold
@@
Query time (for promoted data). Here, we report the query time improvement by CSV for the ‘promoted keys’ (i.e., the keys that is promoted to an upper level of the index by CSV), compared to the original index. We depict the total time saved due to the method in Fig. 6. The general trend is that adding more virtual points (i.e., increasing the smoothing budget, $\alpha$) saves more query times. LIPP and SALI tend to perform quite similarly due to SALI using LIPP as the base index. For LIPP and SALI indexes, the easy to learn datasets (Facebook and Covid) stabilise after a certain number of virtual points are inserted. This is because the original datasets’ CDFs are already quite linear. The same pattern was not observed for ALEX, this is because ALEX has an additional leaf-node search step not required by LIPP and SALI (CSV forms larger nodes that could lead to longer leaf-node search times).
Fig. 7 further reports the average query improvements as a percentage against the original index structure over the promoted data. It shows that applying CSV yields a query time improvement of up to 34%, with stronger benefits observed over the two SOTA index structures LIPP and SALI. Smaller performance gain is observed over ALEX due to its leaf-node search process. It is important to note that CSV still yields consistent query time improvements in this case. As $\alpha$ increases, the relative query time improvements for ALEX also increases, as the leaf-node search efficiency is improved due to increased accuracy of the refitted indexing functions. Since there is no such searching in LIPP and SALI’s query process, their query performance is stagnant and the improvement represents the reduction in the index traversal time for query processing.
(a) Percentage of promoted
data (LIPP)
@@
model reduction (LIPP)
(d) Percentage of promoted
data (SALI)
(e) Percentage of storage
space increase (SALI)
(f) Percentage of index
model reduction (SALI)
(g) Percentage of promoted
data (ALEX)
(h) Percentage of storage
space increase (ALEX)
(i) Percentage of index
model reduction (ALEX)
Figure 8. Space cost vs. smoothing threshold $\alpha$
@@
Size of the promoted data. Figs. 8(a)8(d), and 8(g) show the percentage of keys promoted. For the Facebook dataset, CSV can promote around 60% of all promotable data (i.e., keys at level 3 or below of the original index), while for the Covid dataset, CSV promotes around 30% of the promotable data. For the harder to learn datasets, OSM and Genome, CSV also manages to promote up to 27% and 57% of the promotable data, respectively. The datasets with the most promoted data are again different for ALEX, due to its structural difference. Even though OSM and Genome report lower percentages of promoted data, the actual number of promoted keys is higher. This is because they have much more promotable data (i.e., more keys in lower levels). Overall, as the smoothing threshold increases, more keys get promoted to upper levels. This is consistent with the theoretical analysis as adding more virtual points would allow more keys to be placed in nodes in upper levels. The difficult to learn key sets (OSM and Genome) demonstrate this property the most. This can also be matched with Fig. 6, where the total time saved for OSM and Genome is higher due to higher number of data being promoted for those datasets, compared to the other datasets.
@@
Index size. Due to the addition of virtual points, we expect the storage space consumption to increase. This is reflected in Figs. 8(b)8(e), and 8(h). In most cases, less than 10% of additional storage space is required by the CSV-enhanced indexes compared to the original structures, and in the worst case, less than 31%. The space cost overhead is proportional to $\alpha$, which is also intuitive.
@@
The storage space increase is balanced by the removal of unnecessary nodes, whose data is promoted to higher levels. Figs. 8(c)8(f), and 8(i) report the node reductions achieved by CSV, which are given as the percentage of nodes that are removed relative to the number of nodes at levels 3 or lower of the original indexes. They follow similar patterns to the percentage of promoted data as expected.
@@
Pre-processing time for CSV. The times taken to run CSV to optimise the learned index structures are summarised in Tables 3 and 4 for LIPP and ALEX. As SALI is based on LIPP, where CSV reports a similar performance, we omit its results for brevity.
CSV takes more time to run as $\alpha$ grows, which is consistent with our time complexity analysis. The algorithm running times vary across different datasets, again because the datasets have different difficulties in index learning. While the running times of CSV may seem quite large under certain settings, these are one-off pre-processing costs that can be amortised by the improvements in query time. One may further mitigate the impact of the extra construction times by using the original indexes for queries while constructing a parallel index structure with CSV. Once the CSV-optimised structure is ready, it is switched on for query processing.
Table 3. CSV Pre-processing Time (s) for LIPP
$\alpha$ 0.05 0.1 0.2 0.4 0.8
@@
Table 4. CSV Pre-processing Time (s) for ALEX
$\alpha$ 0.05 0.1 0.2 0.4 0.8
@@
(b) SALI
(c) ALEX
Figure 9. Total time saved vs. dataset cardinality
6.2.2. Impact of Dataset Cardinality.
To demonstrate the scalability of CSV against the dataset cardinality, we repeat the experiments on datasets of 12.5 million to 200 million data points. Fig. 9 shows the total query times saved by applying CSV. For all datasets, the times saved grow with the dataset cardinality, with faster growth being observed on the easier datasets (Facebook and Covid) grows faster. This is because there are not many keys in the lower levels for these datasets when the dataset cardinality is small. These results confirm the scalability of CSV towards dataset cardinality.
(a) Total query time saved (LIPP)
(b) Storage increase (LIPP)
(c) Insertion time increase (LIPP)
(d) Total query time saved (ALEX)
(e) Storage increase (ALEX)
(f) Insertion time increase (ALEX)
Figure 10. Performance results vs. data insertions
6.3. Results on Read-write Workloads
Due to the similar trends between LIPP and SALI indexes, SALI is omitted from the results below for brevity. The default smoothing threshold of 0.1 was used for these experiments.
Query time (promoted data). Figs. 10(a) and 10(d) show the total query times saved by CSV for LIPP and ALEX, respectively, compared to the original index structures, as more batches of data are inserted (recall that each batch consists of $0.1n$ data points). Here, the query times saved are decreasing slightly as more data points are inserted for LIPP, because the inserted data points have a higher chance of colliding with the promoted data points which are now in the upper levels, compared to when they are in lower levels as in the original index structure. For ALEX, the trend is quite similar except for on the OSM dataset, where there are two drops after one and three insertion batches (i.e., $0.1n$ and $0.3n$ data points are inserted). This is because the original index structure’s query times happen to be slightly lower in these two cases.